OpenStudy (anonymous):

Evaluate lim(1/x - 1/tanx) when x approach 0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(\frac{ 1 }{ x }-\frac{ 1 }{ \tan x})\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(\frac{ \tan x -x }{ x \tan x })\] So L'Hopital rule is applied. where i should differentiate both numerator and denominator

zepdrix (zepdrix):

Yah that sounds right. Since you have the form 0/0, you're allowed to apply L'Hop.

zepdrix (zepdrix):

Hmm get's messy after that huh? Hmmm.

OpenStudy (anonymous):

Yeah,its bcome complicated after differentiate.It bcome \[\lim_{x \rightarrow 0}(\frac{ x ^{2}\sec ^{2}x-\tan ^{2}x }{ x ^{2}\tan ^{2}x })\]

OpenStudy (anonymous):

I guess